# Generating a Contact Matrix for Aged Care Settings in Australia: an agent-based model study

### v1

@@ Abstract
This study presents an agent-based model (ABM) developed to simulate staff and resident interactions within a synthetic aged care facility, capturing movement, task execution, and proximity-based contact events across three staff shifts and varying levels of resident care. Contacts were defined by spatial thresholds (1.5 m and 3 m) and cumulative duration, enabling the generation of detailed contact matrices. Simulation results showed that low and medium care residents experienced the highest frequency of interactions, particularly with staff on morning and afternoon shifts, while high care residents and night staff had substantially fewer contacts. Contact rates varied significantly by care level and shift, confirmed through Poisson-based regression modelling. Temporal analyses revealed clustering of high-risk contacts during structured daily routines, especially communal and care activities. An integrated airborne transmission module, seeded with a single infectious staff member, demonstrated that infection risk was highest during high-contact shifts and among medium care residents. Vaccination scenarios reduced predicted transmission by up to 68%, with the greatest impact observed when both staff and residents were vaccinated. These findings highlight the importance of accounting for contact heterogeneity in aged care and demonstrate the utility of ABMs for evaluating targeted infection control strategies in high-risk, enclosed environments.
@@ 1 Introduction
Aged care homes, while essential for providing comprehensive support and care to elderly individuals, are particularly vulnerable to infectious disease outbreaks due to several interrelated factors (Kimball2020Asymptomatic; lorig2021abmreview). Residents in these facilities often have compromised immune systems, which heightens their susceptibility to infections. The close living quarters and communal spaces further exacerbate this risk, creating conditions that facilitate the rapid transmission of pathogens. Moreover, staff members, who have frequent, direct interactions with residents and are regularly exposed to external environments, can inadvertently introduce and propagate infections within the facility. This risk is compounded by the continuous influx of healthcare professionals, visitors, and external service providers, all of whom can act as vectors for disease transmission. Although infection control measures are implemented to mitigate these risks, the rapid evolution of viruses can undermine these efforts, leading to more frequent outbreaks.
The dynamic nature of viral pathogens, characterised by their ability to evolve swiftly, poses a significant challenge to the effectiveness of existing infection control measures. This highlights the importance of understanding and predicting transmission dynamics within aged care settings to mitigate the associated risks. Central to these efforts is the use of contact matrices, which quantify interactions between individuals within a population. The accuracy of infectious disease models, particularly in forecasting outbreaks, is heavily dependent on the quality and specificity of these contact matrices.
One of the most widely used contact matrices, Polymod, was developed through survey-based data collection, where participants recorded details about their interactions over a single day (mossong2008social). This method has been effective in providing broad, population-level estimates of contact patterns, with country-wide matrices regularly updated from more recent surveys. However, Polymod is primarily designed for general population studies, averaging contacts across large age groups, up to 75+, and does not account for the unique contact patterns in specific environments like aged care facilities. In aged care settings, where the interactions are highly specific and vary greatly from the broader population, Polymod’s broad averaging fails to capture the nuanced contact dynamics that are crucial for accurate modelling in these environments.
In response to the limitations of broad-scale contact surveys, several studies have employed agent-based models to generate context-specific contact matrices for use in infectious disease modelling (Baek2021Correction; cuevas2020abmfacilities; Willem2015Optimizing). This approach has been used in workplace settings, schools, and urban communities to simulate heterogeneous interaction structures and inform intervention planning, particularly where empirical contact data are unavailable or infeasible to collect. These simulation-based matrices have also been integrated into compartmental epidemic models to improve realism in transmission estimates (Willem2015Optimizing). However, despite their growing application, relatively few models have focused on aged care environments, where interaction patterns are shaped by care level, staff workflows, and architectural layout.
In such specialised environments, simulation-based methods are often employed to generate contact matrices. However, these methods are typically highly specific to individual locations and lack the generalisability needed for broader applications (lorig2021abmreview; Baek2021Correction; cuevas2020abmfacilities; Willem2015Optimizing). Agent-Based Modelling (ABM) offers a sophisticated and flexible approach to simulate the detailed, location-specific interactions among individuals in aged care environments. ABMs provide a framework to capture the complexities of social, spatial, and temporal dimensions, allowing for a nuanced representation of contact dynamics.
While aged care facilities are high-risk environments for infection spread due to complex and heterogeneous contact patterns, ABMs provide an effective framework for addressing these challenges. They are characterised by highly vulnerable populations and structured interactions that vary by care level and staff role. ABMs capture this heterogeneity by representing individual behaviours, health status, and interaction patterns that drive transmission risk. In addition, they enable the modelling of intricate contact networks, accounting for varying frequencies and durations of interactions, which are crucial for predicting infection dynamics. This study develops a transferable modelling approach to generate high-resolution contact matrices for aged care, evaluating how structured routines and resident–staff interactions shape contact formation and infection risk. The study also explores how changes in proximity thresholds and temporal resolution affect contact frequency and duration, to inform the design of realistic transmission models tailored to high-risk, enclosed settings. By examining how proximity thresholds and temporal resolution influence contact estimates, the model supports the design of context-sensitive transmission frameworks suited to high-risk, enclosed settings.
@@ 2 Methods
In this study an agent-based simulation is used to estimate contact matrices within a synthetic aged care facility parametrised from empirical data collected in a daily activity survey created for this thesis. The agent-based simulation model uses the Anylogic 8 software to analyse interactions and contact patterns within an aged care facility and is designed to simulate the daily activities of residents and staff, capturing the dynamics of movement, task execution, and social interactions.
@@ 2.1 Overview
The model represents the aged care facility as a continuous environment, structured into various functional areas such as bedrooms, bathrooms, dining rooms, and communal spaces. The agents in the model are designed with characteristics, features and behaviours relevant to their roles, such as mobility, care level, and shift schedules. The simulation operates over a continuous time scale with repeating schedule every 24 hours, with the model tracking the location, movement, and interactions of all agents throughout the simulation period. As time in simulation progresses, interactions between agents are recorded as contacts, which are defined based on proximity and duration criteria (see below). These contacts are opportunities for disease transmission and can then be used in further disease modelling such as deterministic or stochastic models at larger scale. The model’s logic includes the generation, prioritisation, and execution of tasks, which are governed by statecharts, task generation functions, and a pedestrian model that manages agent movement within the environment.
The model was developed and is reported using the ODD (Overview, Design concepts, Details) protocol, which provides a standardised approach for describing agent-based models (Grimm2010ODD). The ODD protocol ensures that the model’s structure and processes are clearly defined and reproducible (Grimm2010ODD). Additionally, human decision-making processes within the model were described following the ODD+D extension, which provides a framework for integrating decision-making behaviours in agent-based models [8]. The aim is to create a model which incorporates the behavioural models, task prioritisation, and social force algorithms to reflect the complexities of a real-world aged care environment. Figure 1 provides an overview of the overall model architecture.
@@
2.2 Description
2.2.1 Parametrisation
The behaviours, timings and key parameters governing the day-to-day activities of residents and staff agents in the model was informed through the collection of diarised activity data from aged care workers. This data was also used to develop the suite of behaviours represented within the ABM for the main agent types. The aim of this data collection was to support the development of representative parameters closely mirroring what occurs in real aged care facilities. This data is not available in the peer reviewed literature or publicly available datasets which are often focused on broader health contacts or interventions. The survey comprised of 21 questions of aged care workers informed the model’s parameters, providing insights into regular activities and the frequency and duration of interactions between staff and residents. This information guided the structuring of shift schedules and staff movement patterns within the model. By understanding the frequency and duration of these interactions, the model was parameterised to mirror real-world dynamics. This detailed approach is essential for simulating the nuanced patterns of contact that drive disease transmission in aged care settings. Ethics approval from The University of New South Wales Human Research Ethics Committee (approval number: HC230501) and the aged care provider’s Research and Governance Office approval (Project code: R277). Survey design, data categorisation and details on informing parameters can be found within Annex A. Task duration and time estimates, informed by the survey are provided in Table 1.
@@ 2.2.2 Contact Definition
A contact is defined as an interaction where individuals are within either 1.5 or 3 metres of each other for 3 seconds or more, regardless of whether the contact is physical or non-physical. A contact that exceeds 3 seconds and is defined as a prolonged single contact. The model includes two primary categories of agents: i) residents, and ii) staff. The residents are further subcategorised on their mobility and care level needed.
@@ 2.2.3 Populations
The population consists of staff and resident agents, each initialised with attributes relevant to their roles within the simulation. Staff members are assigned to shifts, with their activities structured around the operational needs of the facility. The initialisation process includes assigning residents to specific wings of the facility and organising staff into shifts. These assignments are essential for structuring the flow of activities and interactions within the simulation, ensuring that the model accurately reflects the dynamics of an aged care environment.
The resident agents are categorised into care levels. The three subcategories are defined as (i) high care, which are fully non-mobile to who have very limited mobility and require substantial or event total assistance (e.g. profound dementia residents); (ii) medium care, which are residents who have progressed into dementia but not fully immobile and have a moderate degree of independence requiring some assistance; and (iii) low care, which are residents who are very early in their onset and are fully mobile and broadly independent.
In the model, staff agents play a vital role in maintaining the daily operations of the aged care facility, providing care and support to residents across all care levels, high care, medium care, and low care. Unlike residents, staff agents are not divided into categories but instead have a responsible wing, which has varying resident care levels. Staff agents are scheduled to work in shifts, ensuring continuous care throughout the day and night. The simulation models their movement within the facility, interactions with residents, and execution of tasks, reflecting the real-world dynamics of caregiving in an aged care environment. This approach ensures that the needs of residents at all care levels are met efficiently and effectively, contributing to the overall well-being of the facility’s population.
@@ 2.2.4 Simulation of Physical Environment
The simulation environment models a single-story aged care home, which represents one of the three broad categories of aged care facility layouts, single-storey structures, multi-storey buildings, and clusters of detached cottages. This single-story design was selected based on data from a representative aged care home, chosen to reflect typical conditions found within the sector. The layout supports the simulation of horizontal movement and interaction patterns, which are essential for understanding the spatial dynamics within such an environment. This approach ensures that the model accurately represents the physical environment in which the agents operate, providing a reliable foundation for analysing resident and staff interactions within the facility.
The facility’s layout is parameterised over an area of approximately 500x500 meters, with specific nodes representing essential locations such as resident bedrooms, bathrooms, dining areas, communal spaces, and staff rooms. These nodes serve as reference points within the simulation, facilitating the movement and interaction of agents throughout the facility.
@@ 2.2.5 Initialisation
The simulation is initialised by configuring the environment to more accurately represent the physical layout of the aged care facility, with functional spaces such as bedrooms, bathrooms, dining rooms, and activity areas. Residents and staff are assigned specific locations based on their care levels and shift schedules, aligning with the real-world operations of the facility. Key operational parameters, including resident numbers and shift timings, are initialised to reflect the facility’s daily dynamics as outlined in Table 2, establishing the foundational conditions under which the model operates. During this initiation phase, decision rules and algorithms are activated to define agent behaviours, such as task prioritisation and interaction patterns. These rules, informed by the survey data and real-world parameters, guide the agents’ decisions throughout the simulation, ensuring that their actions align as closely as possible with real world operational dynamics of the aged care facility. Feedback mechanisms are also initialised to monitor the performance of agents and adjust parameters as needed, ensuring that the model remains responsive to changes in the simulated environment. This includes the adaptation of shift schedules, reallocation of resources, and modification of agent interactions based on the evolving conditions within the simulation.
@@
Table 2: Global parameters within the main agent within the simulation.
@@ 2.2.6 Agent Activities and Scheduling
The activity and schedule management within the simulation is structured around a framework that dynamically governs both staff and resident activities, ensuring that all tasks are systematically generated, prioritised, executed, and completed. The system relies on a combination of statecharts, time-based task generation functions, activity tokens, process models to manage task generation and execution, incorporated into a pedestrian model to simulate physical interactions. These components work together to allow for complex interactions to emerge by implementing certain degrees of individual autonomy, accounting for relevant interactions and the environment within the simulation, ensuring that each agent’s actions align with their roles and the overall objectives of the system. Figure 2 outlines the task management framework for the agents within the model.
Activities within the simulation are initiated and managed through statecharts, which are tailored to each agent type. Each agent is equipped with multiple statecharts that operate simultaneously, each containing conditional information regarding the timing and duration of various behaviours and tasks. These statecharts regulate the flow of activities by determining when an agent should start a task, continue with an ongoing task, or transition to a different state based on evolving conditions within the simulation. Transitions within these statecharts are activated by specific conditions or events, such as the availability of a new task, the completion of a current task, or external triggers. These triggers may be time-based, message-based, or condition-based, and the model continuously monitors for them to decide when a new activity should begin. When a trigger condition is met, the statechart transitions the agent from an idle or waiting state to an active state and triggers the generation of the activity function.

### v2

@@ Abstract
Understanding infectious disease transmission in institutional settings requires modelling approaches that can represent how contacts arise from structured routines, roles, and spatial constraints. In aged care facilities, interactions are shaped by care delivery processes, staff scheduling, and resident mobility, producing contact patterns that differ fundamentally from those assumed in population-level models. However, standard contact matrices used in epidemiological modelling are typically derived from general population surveys and do not capture these institutional mechanisms.
This study develops an agent-based modelling framework to generate high-resolution contact matrices for aged care facilities by simulating task-driven behaviour, staff workflows, and movement through shared spaces. Rather than prescribing contact structure, interactions emerge endogenously from scheduled activities and proximity during task execution. The model is parameterised using collected activity-diary data from aged care workers and is implemented with behavioural logic decoupled from the physical layout, allowing adaptation to alternative facility designs without modifying core mechanisms.
Simulation results show pronounced heterogeneity in contact patterns across resident care levels and staff shifts. Low and medium care residents exhibited substantially higher contact frequencies than high care residents, while staff working day and afternoon shifts accounted for the majority of resident–staff interactions. Temporal analyses revealed clustering of contacts around structured daily routines, including meals and communal activities. Integrating a proximity-based airborne transmission component parameterised for SARS-CoV-2 demonstrated that transmission risk was concentrated during high-contact shifts and among more mobile resident groups. Vaccination scenarios reduced predicted transmission substantially, with the greatest reductions observed when both staff and residents were vaccinated.
By explicitly linking organisational processes to emergent contact structure, this framework provides a reproducible and transferable approach to contact matrix generation for institutional environments. The model supports more realistic transmission modelling and offers a basis for evaluating targeted infection control strategies in high-risk care settings.
@@ 1 Introduction
Infectious disease transmission within institutional settings presents distinct modelling challenges that are not adequately addressed by approaches developed for general population dynamics [21]. In environments such as aged care facilities, interactions are structured by formal roles, scheduled routines, and shared physical spaces, producing contact patterns that differ fundamentally from those assumed in many epidemic models [23, 13]. Understanding how these structured interactions give rise to transmission risk requires modelling frameworks that can represent not only who interacts with whom, but how and why those interactions occur.
Aged care facilities are a particularly pertinent example of this challenge. Daily operations are organised around care delivery, staff allocation, and resident needs, resulting in repeated and prolonged interactions between specific individuals [19, 31, 30, 22, 28, 20]. These interaction patterns are highly heterogeneous, vary across time and space, and are closely tied to organisational processes rather than individual choice alone. As a result, transmission risk in these settings is shaped by the sequencing of activities, staff workflows, and proximity during task execution, rather than by homogeneous mixing within broad population categories [36].
Many infectious disease models rely on contact matrices to summarise interaction patterns between population groups. While effective for large-scale modelling, such representations typically treat contacts as static, externally specified inputs and abstract away the mechanisms that generate them [25]. In institutional environments, this abstraction risks obscuring key drivers of transmission and limiting the interpretability of model outputs [38]. Addressing this limitation requires approaches that can generate contact patterns endogenously from behavioural and organisational processes.
This study develops an agent-based modelling framework to generate contact matrices for aged care facilities in which interactions emerge from routine activities, staff allocation, and movement through shared spaces. Rather than prescribing contact structure in advance, contacts arise from the sequencing of tasks and spatial proximity during care delivery. Behavioural scheduling and task logic are implemented independently of the facility layout, allowing alternative floorplans to be introduced without modifying the underlying interaction mechanisms. The following section situates this approach within existing modelling paradigms and examines common assumptions about contact representation in institutional settings.
2 Background
Mathematical and computational models of infectious disease transmission rely critically on how interpersonal contacts are represented. In many modelling frameworks, contacts are summarised using contact matrices that encode the expected number of interactions between predefined population groups over a given time period [25]. These matrices are typically treated as fixed inputs, specified exogenously to the model dynamics, and are often derived from self-reported contact surveys conducted over short observation windows. Such representations have proven useful for large-scale population modelling, where fine-grained behavioural detail is neither available nor computationally tractable [36].
However, contact matrices encode strong and often implicit assumptions about how interactions arise. By construction, they abstract away the processes through which contacts are generated, collapsing heterogeneous individual behaviours into aggregate summaries. Temporal ordering, task sequencing, and spatial constraints are typically omitted, and repeated interactions between the same individuals are treated equivalently to independent encounters [38]. These assumptions imply that contact opportunities are weakly structured in time and largely independent of organisational context. While such simplifications may be defensible for modelling diffuse community transmission, they become problematic in institutional environments where interactions are shaped by formal roles, routines, and shared spaces [21]. Aged care facilities exemplify such environments. Contacts within these settings do not arise randomly, nor are they evenly distributed across individuals. Instead, interactions are generated through care routines, staff assignment, task prioritisation, and the physical organisation of space [9]. Residents’ contact opportunities depend on care level and mobility, while staff contacts are structured by shift schedules, workload, and responsibility for specific residents or wards. These processes give rise to repeated, temporally clustered interactions between particular individuals, producing exposure patterns that cannot be adequately captured by static or aggregate representations. As a result, models that rely on externally specified contact matrices may mischaracterise both the timing and intensity of transmission opportunities, with implications for estimated outbreak dynamics and intervention effectiveness [3].
Empirical contact matrices derived from general population surveys, such as POLYMOD, illustrate these limitations. While widely used, these matrices are based on short recall periods and broad population categories, and are not designed to represent task-driven or institutionally constrained interactions [25]. Applying such matrices to aged care settings implicitly assumes that contacts within facilities resemble those of the broader community, differing only in scale rather than structure. This assumption obscures key mechanisms of transmission, including prolonged staff–resident interactions, repeated care encounters, and spatial clustering within shared indoor environments [19].
Agent-based models (ABMs) offer an alternative approach by representing individuals as autonomous agents whose interactions emerge from behavioural rules, task execution, and movement through space [13, 12]. Rather than treating contacts as exogenous inputs, ABMs generate contact patterns endogenously as outcomes of simulated social and organisational processes. This generative perspective allows contact formation to be coupled directly to activity schedules, role-based responsibilities, and spatial constraints, making ABMs particularly well suited to modelling institutional settings [21]. Previous studies have used ABMs to generate context-specific contact patterns in schools, workplaces, hospitals, and urban environments, demonstrating that heterogeneity in contact structure can substantially alter predicted transmission dynamics [29, 18]. Despite these advantages, ABM-based contact generation approaches often introduce their own limitations. Many models are tightly coupled to specific physical layouts, datasets, or operational assumptions, limiting their transferability and reproducibility [3]. Behavioural logic is frequently embedded within spatial representations, making it difficult to adapt models to alternative architectural configurations without extensive redevelopment. This coupling constrains the ability to explore how changes in facility design or organisation might influence contact patterns and transmission risk.
From a modelling perspective, this tension highlights a broader trade-off between contextual realism and generalisability. Capturing the mechanisms that generate contacts in institutional environments requires explicit representation of tasks, schedules, and space, yet overly bespoke implementations risk producing models that are informative only for a single setting [21]. Addressing this challenge requires modelling frameworks that treat contact formation as a dynamic outcome of behavioural and organisational processes, while maintaining sufficient modularity to support reuse and adaptation across facilities.
The present study adopts this perspective by conceptualising contacts not as predefined inputs, but as emergent properties of task-driven activity within an institutional environment. Behavioural and scheduling processes are implemented independently of the physical layout, allowing alternative floorplans to be incorporated without altering the core logic governing interaction formation. This design choice reflects a commitment to representing the mechanisms that generate contacts, rather than prescribing their aggregate structure, and provides a foundation for reproducible, institution-specific contact matrix generation.
@@ 3 Methods
In this study, an agent-based simulation is used to estimate contact matrices within a synthetic aged care facility parameterised from empirical data collected in a daily activity survey created for this study. The agent-based simulation model uses the Anylogic 8 software to analyse interactions and contact patterns within an aged care facility, and is designed to simulate the daily activities of residents and staff, capturing the dynamics of movement, task execution, and social interactions.
The model was developed and documented following the ODD (Overview, Design concepts, Details) protocol for agent-based models [14]. Human decision-making processes were additionally specified using the ODD+D extension, which supports the explicit representation of behavioural decision-making in agent-based systems [26]. The model integrates behavioural rules, task prioritisation, and social force algorithms to reflect the operational complexities of a real-world aged care environment.
@@ 3.1 Overview
The model represents the aged care facility as a continuous environment, structured into various functional areas such as bedrooms, bathrooms, dining rooms, and communal spaces. The agents in the model are designed with characteristics, features and behaviours relevant to their roles, such as mobility, care level, and shift schedules. The simulation operates over a continuous time scale with repeating schedule every 24 hours, with the model tracking the location, movement, and interactions of all agents throughout the simulation period. As time in simulation progresses, interactions between agents are recorded as contacts, which are defined based on proximity and duration criteria. These contacts are opportunities for disease transmission and can then be used in further disease modelling such as deterministic or stochastic models at larger scale. The model’s logic includes the generation, prioritisation, and execution of tasks, which are governed by statecharts, task generation functions, and a pedestrian model that manages agent movement within the environment [Figure 1].
@@
3.2 Design Concepts
3.2.1 Parameterisation
The behaviours, timings and key parameters governing the day-to-day activities of residents and staff agents in the model was informed through the collection of diarised activity data from aged care workers. This data was also used to develop the suite of behaviours represented within the ABM for the main agent types. The aim of this data collection was to support the development of representative parameters closely mirroring what occurs in real aged care facilities. This data is not available in the peer reviewed literature or publicly available datasets, which are often focused on broader health contacts or interventions. The survey comprised of 21 questions of aged care workers informed the model’s parameters, providing insights into regular activities and the frequency and duration of interactions between staff and residents. This information guided the structuring of shift schedules and staff movement patterns within the model. By understanding the frequency and duration of these interactions, the model was parameterised to mirror real-world dynamics. This detailed approach is essential for simulating the nuanced patterns of contact that drive disease transmission in aged care settings.
@@ 3.2.2 Contact Definition
A contact is defined as an interaction in which two agents remain within either 1.5 or 3 metres of each other for duration of at least 3 seconds or more, regardless of whether the contact is physical or non-physical. This duration threshold was used to exclude transient co-presence events that are unlikely to result in meaningful exposure. Contacts are recorded as single continuous events and persist across consecutive simulation time-steps until the agents separate beyond the specified distance threshold.
@@ 3.2.3 Populations
The model includes two primary categories of agents: i) residents, and ii) staff. The residents are further subcategorised on their mobility and care level needed. The population consists of staff and resident agents, each initialised with attributes relevant to their roles within the simulation. Staff members are assigned to shifts, with their activities structured around the operational needs of the facility. The initialisation process includes assigning residents to specific wings of the facility and organising staff into shifts. These assignments are essential for structuring the flow of activities and interactions within the simulation, ensuring that the model accurately reflects the dynamics of an aged care environment.
The resident agents are categorised into care levels. The three subcategories are defined as (i) high care, which are fully non-mobile to who have very limited mobility and require substantial or event total assistance, such as progressed dementia patients; (ii) medium care, which are residents who have progressed into dementia but not fully immobile and have a moderate degree of independence requiring some assistance; and (iii) low care, which are residents who are very early in their onset and are fully mobile and broadly independent.
In the model, staff agents play a vital role in maintaining the daily operations of the aged care facility, providing care and support to residents across all care levels, high care, medium care, and low care. Unlike residents, staff agents are not divided into categories, but instead have a responsible wing, which has varying resident care levels. Staff agents are scheduled to work in shifts, ensuring continuous care throughout the day and night. The simulation models their movement within the facility, interactions with residents, and execution of tasks, reflecting the real-world dynamics of caregiving in an aged care environment. This approach ensures that the needs of residents at all care levels are met efficiently and effectively, contributing to the overall well-being of the facility’s population.
3.3 Details
@@ 3.3.1 Simulation of Physical Environment
The simulation environment represents a single-storey aged care facility, parameterised from a representative real-world layout to support realistic movement and interaction patterns. The facility’s layout is parameterised over an area of approximately 500x500 meters, with specific nodes representing essential locations such as resident bedrooms, bathrooms, dining areas, communal spaces, and staff rooms. These nodes serve as reference points within the simulation, facilitating the movement and interaction of agents throughout the facility.
The spatial layout is implemented as a standalone model component, allowing alternative floorplans to be incorporated by modifying only the layout structure, while agent behaviour, scheduling logic, and contact-generation processes remain unchanged.
@@ 3.3.2 Initialisation
The simulation is initialised by configuring the environment to more accurately represent the physical layout of the aged care facility, with functional spaces such as bedrooms, bathrooms, dining rooms, and activity areas. Residents and staff are assigned specific locations based on their care levels and shift schedules, aligning with the real-world operations of the facility. Key operational parameters, including resident numbers and shift timings, are initialised to reflect the facility’s daily dynamics as outlined in Table 2, establishing the foundational conditions under which the model operates. During this initiation phase, decision rules and algorithms are activated to define agent behaviours, such as task prioritisation and interaction patterns. These rules, informed by the survey data and real-world parameters, guide the agents’ decisions throughout the simulation, ensuring that their actions align as closely as possible with real world operational dynamics of the aged care facility. During initialisation, behavioural rules and task priorities are activated to govern agent actions throughout the simulation.
@@
Table 2: Model-level parameters fixed at initialisation and applied globally throughout the simulation.
@@ 3.3.3 Agent Activities and Scheduling
The activity and schedule management within the simulation is structured around a framework that dynamically governs both staff and resident activities, ensuring that all tasks are systematically generated, prioritised, executed, and completed. The system relies on a combination of statecharts, time-based task generation functions, activity tokens, process models to manage task generation and execution, incorporated into a pedestrian model to simulate physical interactions. These components work together to allow for complex interactions to emerge by implementing certain degrees of individual autonomy, accounting for relevant interactions and the environment within the simulation, ensuring that each agent’s actions align with their roles and the overall objectives of the system. The activity workflow combines a process model, which initiates and schedules tasks, with a statechart that governs the agent’s active task execution [Figure 2].
Activities within the simulation are initiated and managed through statecharts, which are tailored to each agent type. Each agent is equipped with multiple statecharts that operate simultaneously, each containing conditional information regarding the timing and duration of various behaviours and tasks. These statecharts regulate the flow of activities by determining when an agent should start a task, continue with an ongoing task, or transition to a different state based on evolving conditions within the simulation. Transitions within these statecharts are activated by specific conditions or events, such as the availability of a new task, the completion of a current task, or external triggers. These triggers, or signals, may be time-based, message-based, or condition-based, and the model continuously monitors for them to decide when a new activity should begin. When a trigger condition is met, the statechart transitions the agent from an idle or waiting state to an active state and triggers the generation of the activity function.
To prevent indefinite queuing of low-priority tasks under sustained workload, the scheduling process includes a queue timeout rule. Tasks that remain queued for more than 40 minutes without execution are either escalated in priority or terminated depending on task type, ensuring bounded waiting times and stable system behaviour.